# Demystifying Language Model Forgetting with Low-rank Example Associations

### v3

@@ Abstract
Large Language models (LLMs) suffer from forgetting of upstream data when fine-tuned. Despite efforts on mitigating forgetting, few have investigated whether, and how forgotten upstream examples are dependent on and associated with newly learned tasks. Insights on such associations enable efficient and targeted mitigation of forgetting. In this paper, we empirically analyze forgetting (measured in log-perplexity increase) that occurs in $N$ upstream examples of language modeling or instruction-tuning after fine-tuning LLMs on one of $M$ new tasks, visualized in $M\times N$ matrices. We demonstrate that the matrices display simple low-rank patterns, often well-approximated with multiplicative scalar effects of upstream examples and newly learned tasks. We also examine fine-grained associations with visualization and statistics. Leveraging the low-rank nature of the associations, we predict forgetting of upstream examples when fine-tuning on unseen tasks with matrix completion over the empirical associations. This enables fast identification of most forgotten examples without expensive inference on the entire upstream data. The approach, despite simplicity, outperforms prior approaches that learn semantic relationships of learned tasks and upstream examples with LMs for predicting forgetting. We demonstrate the practical utility of our analysis by showing statistically significantly reduced forgetting as we upweight predicted examples for replay at fine-tuning.
@@ 1 Introduction
There has been growing need for promptly updating LLMs to mitigate harmful behaviors, update outdated knowledge, and expand the set of application domains. Although fine-tuning allows efficient and incremental updates of models, it risks catastrophic forgetting (McCloskey & Cohen, 1989; Goodfellow et al., 2014) of upstream knowledge learned in the pre-training or instruction-tuning phase, causing unintended prediction changes over known information. This is problematic for the stability of online deployed LLM systems, limiting the feasibility of continual fine-tuning in practice (Raffel, 2023; Shi et al., 2024).
While extensive works have developed algorithms to mitigate forgetting (Wu et al., 2024), the increasing scale of the model and data necessitates more efficient and targeted approaches. A key aspect of the challenge is understanding when and what LLMs forget, prompting analysis into the patterns of frequently forgotten examples (Toneva et al., 2019; Maini et al., 2022; Zhang & Wu, 2024), and the impacts of models and hyperparameters on forgetting (Mirzadeh et al., 2022; Kalajdzievski, 2024; Que et al., 2024; Ibrahim et al., 2024). However, how the associations between learned tasks and upstream examples inform forgetting remains under-explored. Understanding such associations creates potential for targeted mitigation of forgetting over specific upstream examples when a new task is learned. While theoretical and empirical study indicate associations between learned and forgotten tasks in shallower neural networks (Lee et al., 2021; Goldfarb et al., 2024; Ramasesh et al., 2021), the associations are under-explored for LLMs, or measured regarding upstream data of language modeling or instruction-tuning.
In this paper, we empirically study the associations between learned tasks and forgotten upstream examples of language modeling or instruction-tuning. We experiment with OLMo-1B, OLMo-7B, OLMo-7B-Instruct (Groeneveld et al., 2024) and MPT-7B (Team, 2023) models where upstream data is open-source. We analyze forgetting (in log perplexity increase) over $N$ upstream examples, after fine-tuning the model on one of $M$ unseen instruction-tuning tasks, and represent the results in a $M\times N$ matrix. Afterwards, we visualize the matrices and fit the observations with statistical models to analyze the associations. Figure 1 illustrates the analysis setup.
Our findings suggest that the associations between learned tasks and forgotten examples are usually simple and display a low-rank pattern, as simple regression models can decently fit of the associations. We extract more fine-grained associations with visualization and statistics, showing cases where specific upstream examples are forgotten when learning a new task. We demonstrate that forgetting does not correlate with many predefined similarity metrics of learned tasks and upstream examples, such as token overlap or gradient dot-products, making these metrics poor predictors of forgetting. Inspired by the findings, we directly utilize statistics of forgetting to predict example forgetting on unseen tasks by solving a matrix completion problem over the association matrices, analogical to collaborative filtering (Sarwar et al., 2001) in recommender systems, achieving both efficiency and interpretability. Our $k$-nearest neighbor (KNN) model outperforms prior approaches that learn semantic relations of two examples with LMs (Jin & Ren, 2024) . We verify the benefit of prediction by upweighting examples with higher predicted forgetting during replay as we fine-tune LLMs on new instruction-tuning tasks, achieving statistically significant improvement in alleviating forgetting compared to replaying random examples.
@@
Figure 1: The problem setup of analyzing the associations between learned tasks and forgotten upstream examples as we fine-tune LLMs on one of unseen new tasks. Over total $N$ upstream examples and $M$ unseen tasks, we measure and record forgetting in a $M\times N$ matrix and analyze the statistical properties of the associations.
@@
To summarize, the contributions of this paper are (1) an empirical analysis on how forgotten examples are associated with learned tasks in representative 1B and 7B language models, and (2) a novel approach of predicting example forgetting by solving a matrix completion problem over the empirical associations, and (3) a practical and efficient algorithm to mitigate forgetting during LLM fine-tuning by upweighting upstream examples for replay according to the predicted forgetting.
@@ 2 Problem and Analysis Setup
In this section, we start by formally defining the metrics of forgetting and set up the problem formulation of analyzing the associations between learned tasks and forgotten examples. We then introduce models and datasets used for collecting the statistics.
@@ 2.1 Collecting Statistics of Forgetting
Upstream examples and learned tasks. Large Language models (LLMs) are commonly pre-trained with language modeling objectives over a massive collection for corpora, with some LLMs further post-trained (instruction-tuned) to better follow human instructions. We use upstream data to refer to language modeling or instruction tuning training data used at the pre-training or post-training phase of LLMs. For upstream data of language modeling, we define each upstream example $x_{j}\in x_{1..N}$ as a chunk of document (e.g., a Wikipedia article) of a model-specific maximum number of tokens. For instruction tuning, each $x_{j}\in x_{1..N}$ corresponds to a pair of instructions and ground truth responses.
Measuring forgetting. We fine-tune an LLM (or an instruction-tuned LLM) on one unseen instruction-tuning task $T_{i}$ from a collection of tasks $T_{1..M}$. This results in $M$ separately fine-tuned models $f_{1..M}$. We then evaluate performance degradation on each upstream example $x_{i}\in x_{1..N}$. We measure log perplexity as the performance metric as they are applicable to both language modeling and instruction tuning, and are known to correlate well with other dataset-specific metrics (Hoffmann et al., 2022). For language modeling, log perplexity of an example $x_{i}$ is averaged over all tokens in the example; for instruction tuning, we average log perplexity over the ground truth response tokens only. We measure forgetting $z_{ij}$ that occurs on an upstream example $x_{j}\in x_{1..N}$ as increase (degradation) in log perplexity after fine-tuning the LM on a new task $T_{i}\in T_{1..M}$. We record forgetting $z_{ij}$ in an association matrix $Z\in\mathbb{R}^{M\times N}$.
@@ 2.2 Models and Datasets
Our analysis requires access to upstream data of LLMs for measuring their log perplexity changes occurred after fine-tuning. We experiment with OLMo-1B, OLMo-7B, OLMo-7B-Instruct, and MPT-7B where upstream data of language modeling and instruction-tuning is open-source.
OLMo-1B and 7B. OLMo models are pre-trained on Dolma (Soldaini et al., 2024), a massive collection of text covering diverse domains such as Wikipedia and news articles. We fine-tune LMs over 66 tasks from FLAN-V2 (Longpre et al., 2023), 11 tasks from Tulu V2 (Ivison et al., 2023), and 8 tasks from Dolly (Conover et al., 2023), obtaining 85 fine-tuned models. The collections cover diverse instruction-tuning tasks such as reading comprehension, math reasoning, and safety alignment. We then evaluate log perplexity increase on a 1% subset of Dolma-v1.6-Sample. Each upstream example is a maximum 2,048-token document from Dolma, resulting in 141,816 examples. For 7B models, We perform either full-parameter fine-tuning or LoRA adapter tuning (Hu et al., 2022) only, noted as OLMo-7B (Full FT) or OLMo-7B (LoRA) respectively.
OLMo-7B-Instruct. OLMo-7B-Instruct models are instruction-tuned on Tulu V2. In our experiments, we few-shot fine-tune OLMo-7B-Instruct over new task data from MMLU (Hendrycks et al., 2021), BBH (Suzgun et al., 2022), TruthfulQA (Lin et al., 2022), and Dolly, and evaluate log perplexity increase over a stratified sample of 10,718 examples from Tulu v2 as upstream examples. The number of training examples in these datasets are smaller (on the scale of tens to hundreds of examples), making the setup closer to LLM error editing (Yao et al., 2023; Zhang et al., 2024). We perform either full-parameter fine-tuning or LoRA fine-tuning.
MPT-7B. MPT models are pre-trained on a diverse collection of corpora. We fine-tune the LM over the same 85 instruction-tuning tasks as OLMo models and evaluate forgetting on Redpajama (Computer, 2023), involved as a part of the pretraining corpora of MPT. We sample 10,000 maximum 2,048 token documents from Redpajama as upstream examples.
We use a learning rate of $2e^{-6}$ for full-parameter fine-tuning and $1e^{-4}$ for LoRA fine-tuning. We include the other training details in Appendix A.
(a) OLMo-1B; forgetting over Dolma after full-parameter fine-tuning on FLAN and Tulu.
(b) OLMo-7B; forgetting over Dolma after full-parameter fine-tuning on FLAN and Tulu.
(c) OLMo-7B-Instruct; forgetting over Tulu after full-parameter fine-tuning on unseen instruction-tuning tasks.
Figure 2: Visualized matrices of associations between learned tasks and forgotten examples. We plot forgetting (log-perplexity increase) that occurs on an upstream example (in $x$-axis) after learning a new task (in $y$-axis). Log-perplexity increase can be zero or negative, which implies no forgetting.
@@ 3 Associations between Learned Tasks and Forgotten Examples
In this section, we address the research questions regarding the associations between learned tasks and forgotten upstream examples represented in the $M\times N$ association matrices $Z$. We start by examining how complicated are the associations represented by the matrices $Z$ (Sec. 3.1). We then extract and analyze more fine-grained associations in $Z$ (Sec. 3.2), and examine how the associations correlate with similarity measures of learned tasks and upstream examples (Sec.3.3).
3.1 How Complicated are the Associations?
In a hypothetical extreme case, the associations between the learned tasks and forgotten examples can be very simple, e.g., where upstream examples are forgotten regardless of learned tasks; or it can be highly complicated, e.g., where different new tasks cause very specific subsets of upstream examples to be forgotten. Figure 1 (right) illustrates how such simple or complicated associations may display in visualized matrices. To understand the actual associations, we start by visualizing the association matrix $Z$ collected in the setups described in Sec. 2.2, and perform quantitative analysis as we examine whether simple regression models can fit $Z$ with low error.
Analysis on the visualized association matrices. We visualize the association matrices $Z$ in Figure 2 for OLMo-1B, OLMo-7B, OLMo-7B-Instruct as we perform full-parameter fine-tuning. The matrices of MPT and LoRA fine-tuning of OLMo-7B and OLMo-7B-Instruct are presented in Figure 7 in Appendix. We see $Z$ generally displays a neat and simple pattern. We notice that certain upstream are more prone to forgetting, while some are never forgotten (displayed as all-zero columns). Nevertheless, in some cases, upstream examples that are never forgotten elsewhere are forgotten by learning specific new tasks (e.g. wizardlm examples are almost only forgotten after learning tasks from TruthfulQA in Figure 2(c)). It implies the association is a mixture of simple and more complicated patterns.
Quantitative evaluation of simplicity of the associations. We quantitatively measure how well the association matrices $Z$ can be approximated with simple regression models with a small number of learnable parameters. We consider (1) additive linear models, where $z_{ij}=b+\alpha_{i}+\beta_{j}+\epsilon$, where $\alpha_{i}$ and $\beta_{j}$ are learnable parameters associated with each new task or upstream example. (2) multiplicative models (SVD with rank $r$=1), where $z_{ij}=s\alpha_{i}\beta_{j}+\epsilon$. Both models involve $M+N$ learnable parameters plus a bias term. We then measure $R^{2}$ as the metrics of determining how well the regression models fit the association matrices $Z$. Let $f_{ij}$ be the fitted value, $R^{2}$ is defined as $1-\sum_{i,j}(z_{ij}-f_{ij})^{2}/\sum_{i,j}(z_{ij}-\bar{Z})^{2}$. $R^{2}$ ranges between 0 and 1, indicating the portion of total variance in $Z$ that is explained by the regression models. We report $R^{2}$ of additive and multiplicative models under different setups in Figure 3 (a).
Multiplicative models fit the associations better. In 5 out of 6 setups presented in Figure 3 (a), the multiplicative model achieves a better fit than additive models at the same number of trainable parameters. An interpretation of the multiplicative models is that each upstream example is associated with their tendency of being forgotten $(\beta_{j})$; the learned tasks trivially determine how fast all upstream examples are forgotten (with $\alpha_{i}$). This also creates cases where certain upstream are never forgotten $(\beta_{j}\approx 0)$, or learn certain tasks causing little forgetting on any upstream examples $(\alpha_{i}\approx 0)$.
(a) Fitting $Z$ with additive or multiplicative models
(b) Fitting $Z$ with progressively higher rank matrices
Figure 3: $R^{2}$ of fitting the associations between learned tasks and forgotten examples with (a) simple additive linear or multiplicative models, or (b) progressively more expressive models.
A large portion of the variance in the association can be explained by the multiplicative model. We notice the multiplicative models achieve $R^{2}$ between 0.43 to 0.72 in different setups despite their simplicity, suggesting generally simple associations between learned tasks and forgotten examples. This finding is emphasized by the diversity of the learned tasks we considered (spanning from reading comprehension, coding, to safety alignment) and the broad coverage of domains in upstream language modeling and instruction-tuning data. Nevertheless, we also notice the $R^{2}$ scores are relatively lower (0.4 to 0.5) on OLMo-7B and OLMo-7B-Instruct, implying more complicated associations than the other setups (OLMo-1B, MPT-7B, and LoRA fine-tuning of OLMo-7B models), which elicits our next research question.
3.2 What are More Fine-Grained Associations?

### v4

@@ Abstract
Large Language models (LLMs) suffer from forgetting of upstream data when fine-tuned. Despite efforts on mitigating forgetting, few have investigated whether, and how forgotten upstream examples are dependent on newly learned tasks. Insights on such dependencies enable efficient and targeted mitigation of forgetting. In this paper, we empirically analyze forgetting that occurs in $N$ upstream examples of language modeling or instruction-tuning after fine-tuning LLMs on one of $M$ new tasks, visualized in $M\times N$ matrices. We show that the matrices are often well-approximated with low-rank matrices, indicating the dominance of simple associations between the learned tasks and forgotten upstream examples. Leveraging the analysis, we predict forgetting of upstream examples when fine-tuning on unseen tasks with matrix completion over the empirical associations. This enables fast identification of most forgotten examples without expensive inference on the entire upstream data. The approach, despite simplicity, outperforms prior approaches that learn semantic relationships of learned tasks and upstream examples with LMs for predicting forgetting. We demonstrate the practical utility of our analysis by showing statistically significantly reduced forgetting as we upweight predicted examples for replay at fine-tuning.
Catastrophic forgetting; Large language models
@@ 1 Introduction
There has been growing a need for continued fine-tuning of LLMs to mitigate harmful behaviors, update outdated knowledge, and adapt to unseen tasks and domains. Although fine-tuning allows efficient and incremental adaptation of models, models may suffer from catastrophic forgetting (McCloskey & Cohen, 1989; Goodfellow et al., 2014) of upstream knowledge learned in the pre-training or instruction-tuning phase, causing unintended prediction changes over known information. This is problematic for the performance and reliability of online deployed LLMs, limiting the feasibility of continual fine-tuning in practice (Raffel, 2023; Shi et al., 2024).
Existing works demonstrate that replaying or mixing in past examples are effective and scalable approaches to mitigate LLM forgetting (Scialom et al., 2022; Roth et al., 2024; Li et al., 2024a; Ibrahim et al., 2024; Ye et al., 2024). These approaches, however, often rely on random sampling of past examples; knowing what models forget after fine-tuning allows more efficient and targeted mitigation of forgetting – e.g., by prioritizing the replay of more forgotten examples (Toneva et al., 2019; Aljundi et al., 2019a). In this paper, we explore how forgotten upstream examples depend on the learned tasks, so that forgetting caused by unseen tasks can be efficiently inferred from the forgetting that happened while learning other tasks. The complexity of such associations between learned tasks and forgotten examples plays an important role in the predictability; Figure 1 (a) illustrates a hypothetical scenario where certain upstream examples suffer more from forgetting regardless of which tasks are learned, making forgetting easily predictable; in contrast, (b) exemplifies upstream example forgetting that is highly dependent on the learned tasks. Existing theoretical and empirical study on the associations between learned and forgotten tasks focus on shallower models (Lee et al., 2021; Goldfarb et al., 2024; Ramasesh et al., 2021); the problem is under-explored for LLM forgetting or in an example level.  Swayamdipta et al. (2020); Maini et al. (2022) characterize examples that are prone to forgetting, but they do not touch on how example forgetting depends on the learned tasks.
@@
Figure 1: The problem setup of analyzing the associations between learned tasks and forgotten upstream examples as we fine-tune LLMs on one of unseen new tasks. Over total $N$ upstream examples and $M$ unseen tasks, we measure and record forgetting (in red) in a $M\times N$ matrix and attempt to fit the associations with low-rank approximations. Better approximations of low-rank approximations indicate simpler associations between learned tasks and forgotten upstream examples.
Specifically, we start with analyzing the associations between the learned tasks and forgotten upstream examples in LLM fine-tuning. We measure forgetting (in continuous log perplexity increase or binary exact match drop) over $N$ upstream examples, after fine-tuning the model on one of $M$ unseen instruction-tuning tasks, while summarizing the results in a $M\times N$ matrix. We evaluate the complexity of the associations by measuring the goodness-of-fit of low-rank approximations of the example associations. We then examine how the complexity of the example associations varies across model types (OLMo, OLMo2, MPT, Pythia) and sizes (1B to 13B parameters).
Our findings suggest that the associations between learned tasks and forgotten examples are often well-approximated with low-rank matrices. On OLMo-1B and OLMo-7B, rank-3 approximation fits the associations between 85 learned tasks and 140,000 upstream examples with $R^{2}>0.69$. We notice that the forgetting of more capable and recent LLM families are more complicated, requiring higher-rank approximations; within the same model family, the complexity of the associations remains stable or increases with the model size. The matrix decomposition further interprets the associations by distinguishing forgetting that are independent of or dependent on what the model learns.
Following the low-rank approximations of the associations, we predict example forgetting on unseen tasks by solving a matrix completion problem over the association matrices, analogical to collaborative filtering (Sarwar et al., 2001) in recommender systems, achieving both efficiency and interpretability. Our matrix factorization (MF) or $k$-nearest neighbor (KNN) models outperform prior approaches that learn semantic relations of two examples with LMs (Jin & Ren, 2024). As an example, we achieve 57.35 F1 in predicting binary example forgetting where the F1 of random guess is only 6.3.
Lastly, we demonstrate the practical benefit of predicting forgetting in mitigating forgetting. We upweight upstream examples with higher predicted forgetting during replay as we fine-tune LLMs on new instruction-tuning tasks, achieving statistically significant improvement in alleviating forgetting compared to replaying random examples.
@@
To summarize, the contributions of this paper are (1) an empirical analysis on how forgotten examples are associated with learned tasks in representative 1B to 13B language models, and (2) a novel approach of predicting example forgetting by solving a matrix completion problem over the empirical associations, and (3) a scalable and efficient algorithm to mitigate forgetting during LLM fine-tuning by upweighting upstream examples for replay according to the predicted forgetting.
@@ 2 Problem and Analysis Setup
In this section, we start by formally defining the metrics of forgetting and set up the problem formulation of analyzing the associations between learned tasks and forgotten upstream examples. We then introduce models and datasets used for collecting the statistics.
@@ 2.1 Collecting Statistics of Forgetting
Upstream examples and learned tasks. LLMs are commonly pre-trained with language modeling objectives over a massive collection of corpora, and optionally post-trained (instruction-tuned) to better follow human instructions. We use upstream data to refer to language modeling or instruction tuning training data used at the pre-training or post-training phase of LLMs. For upstream data of language modeling, we define each upstream example $x_{j}\in x_{1..N}$ as a chunk of document (e.g., a Wikipedia article) of a model-specific maximum number of tokens. For instruction tuning, each $x_{j}\in x_{1..N}$ corresponds to a pair of instructions and ground truth responses.
Measuring forgetting. We fine-tune an LLM (or an instruction-tuned LLM) on one unseen instruction-tuning task $T_{i}$ from a collection of tasks $T_{1..M}$. This results in $M$ separately fine-tuned models $f_{1..M}$. We then evaluate performance degradation on each upstream example $x_{i}\in x_{1..N}$. We mainly measure log perplexity as the performance metric as they are applicable to both language modeling and instruction tuning, and are known to correlate well with other dataset-specific metrics (Hoffmann et al., 2022). For instruction tuning tasks with a restricted output space (e.g., multi-choice questions), we also measure binary exact matches (EM). We measure forgetting $z_{ij}$ that occurs on an upstream example $x_{j}\in x_{1..N}$ as increase in log perplexity or drop of exact match after fine-tuning the LM on a new task $T_{i}\in T_{1..M}$. We record forgetting $z_{ij}$ in an association matrix $Z\in\mathbb{R}^{M\times N}$.
(a) OLMo-7B; forgetting (in red) over Dolma after full-parameter fine-tuning on FLAN, Tulu, and Dolly.
(b) OLMo-7B-Instruct; forgetting (in red) over Tulu after full-parameter fine-tuning on MMLU, BBH, TruthfulQA, and Dolly.
Figure 2: Examples of visualized matrices of associations $Z$ between $M$ learned tasks and $N$ upstream examples. We plot forgetting (log-perplexity increase) that occurs on an upstream example (in $x$-axis) after learning a new task (in $y$-axis). Log-perplexity increase can be zero or negative, which implies no forgetting. We visualize binary forgetting measured with exact match drops in Fig. 10 in Appendix.
@@ 2.2 Models and Datasets
Our analysis requires open access to upstream data of LLMs. We mainly experiment with open-data LLMs including OLMo-1B, OLMo-7B, and OLMo-7B-Instruct (Groeneveld et al., 2024). We also include OLMo2-7B, OLMo2-13B (OLMo et al., 2024), MPT-7B (Computer, 2023), and Pythia-1B to 12B (Biderman et al., 2023) for studying the complexity of the example associations across model types and sizes.
[…]
Unseen Task $T_{1..M}$ where models are fine-tuned. We fine-tune all non-instruction-tuned models over 66 task from FLAN, 11 tasks from Tulu and 8 tasks from Dolly (Conover et al., 2023). For OLMo-Instruct, we few-shot fine-tune OLMo-7B-Instruct over new task data from MMLU (Hendrycks et al., 2021), BBH (Suzgun et al., 2022), TruthfulQA (Lin et al., 2022), and Dolly. For all models and tasks, we use a learning rate of $2e^{-6}$ for full-parameter fine-tuning and $1e^{-4}$ for LoRA fine-tuning. We include the remaining training details in Appendix A.
@@ 3 Associations between Learned Tasks and Forgotten Examples
(a) Fitting Log PPL increase
(b) Fitting EM Drop
Figure 3: $R^{2}$ and F1 of low-rank approximations of the example associations with progressively higher rank $r$. Forgetting is measured with log perplexity increase or exact match (EM) drop respectively. The number of newly learned tasks (a) $M=85$ (b) $M=124$; upstream examples $N>10,000$ in both cases. “R” notes for random guess of with the positive class probability in (b).
In this section, we analyze the associations between learned tasks and forgotten upstream examples represented in the $M\times N$ association matrices $Z$. We visualize the association matrix $Z$ collected in the setups described in in Figure 2. We formally define low-rank approximations and set up quantitative metrics of the complexity of the associations in Sec. 3.1, and examine the results of approximation across model types and sizes in Sec. 3.2. Lastly, we try to interpret the extracted associations in Sec. 3.3.
3.1 Methods and Metrics of Approximating Example Associations
To examine whether simple patterns are dominant in the example associations represented by $Z$, we attempt to approximate $Z$ with low rank matrices. When $Z$ represents the increase of log perplexity (also the loss of language modeling), we fit matrix factorization models $Z^{r}=\sum_{k=1}^{r}\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$ that minimize the Frobenius norm $||Z-Z^{r}||_{F}$, where $\bm{\alpha}_{k}\in\mathbb{R}^{M}$, $\bm{\beta}_{k}\in\mathbb{R}^{N}$, $r$ is the rank of the matrix decomposition and $k$ is the index of the component. For binary forgetting measured with exact match drop, we fit a logistic matrix factorization model $Z^{r}=\sigma(\sum_{k=1}^{r}\bm{\alpha}_{k}\bm{\beta}_{k}^{T})$ that minimizes the cross entropy between $Z^{r}$ and $Z$, where $\sigma$ is the sigmoid function. We measure $R^{2}$ or F1-scores of the approximation as the goodness-of-fit metrics.
Interpretations. When $r=1$, the approximation effectively assumes that the forgetting $z_{ij}$ (or its logit) is a scalar product of a parameter $\alpha^{(j)}$ specific to each upstream example and each newly learned task $\beta^{(i)}$. The set of more forgotten upstream examples is independent of which task the model learns, as $\beta^{(i)}$ trivially determines how fast forgetting uniformly happens on all upstream examples. With a higher rank $r$, the approximation captures task-dependent forgetting where certain upstream examples are disproportionally more forgotten when learning certain tasks. This inner product formulation is also connected to the first-order approximation of loss increase as an inner product of weight updates and gradients (Lopez-Paz & Ranzato, 2017; Lee et al., 2019), where $r$ is the number of LLM parameters.
3.2 Examining Complexity of Example Associations
(a) $R^{2},r=3$
(b) $R^{2},r=5$
(c) Avg. Forgetting
Figure 4: $R^{2}$ of approximating the example associations $Z$ in $M=19$ Tulu and Dolly tasks and $N=10,000$ randomly sampled upstream examples across different model types and sizes. We present reconstruction with rank $r=3$ (a) or $r=5$ (b), and the average forgetting of models in (c).
General findings. We present $R^{2}$ or F1 of fitting the association with $Z^{r}$ with progressively larger rank $r$ in Figure 3. We notice that across all training setups of OLMo models, $R^{2}$ quickly increases to 0.69 with $r=3$. Notably, even the rank-1 approximation $Z^{1}$ can achieve $R^{2}$ scores higher than 0.5. We see similar findings are on binary forgetting measured with OLMo-Instruct models, where F1 can surpass 0.5 with $Z^{1}$, and 0.6 with $Z^{4}$. The results suggest that simple patterns are dominant in the associations between learned tasks and forgotten examples.
Example associations across model types and sizes. We compare $R^{2}$ of approximations with a fixed $M$ and $N$ over Pythia, MPT, OLMo, and OLMo2 models. Figure 4 summarizes the $R^{2}$ at a given rank $r$. We notice that the goodness of approximations differ among model types. On Pythia and MPT, $R^{2}$ scores at $Z^{3}$ are higher than 0.88, while on OLMo-7B and OLMo2, the $R^{2}$ scores are around 0.75 and 0.65. The size of the models within the same model family also has an impact on $R^{2}$ On Pythia and OLMo2, $R^{2}$ stays stable with slight decrease as the model size increases from 1B to 13B. On OLMo, $R^{2}$ is noticeably lower on 7B models compared to 1B. Model families that forget more (e.g., Pythia) tend to yield simpler example associations (higher $R^{2}$). However, within the same model familiy, OLMo-7B results in a lower $R^{2}$ score forgets than OLMo-1B despite that the average forgetting is higher.
To summarize, we empirically notice that the associations between learned tasks and forgotten upstream examples are more complicated in more recent and capable LLMs, requiring higher-rank approximations. The complexity of the associations stays stable or increases with larger models within the same family. In Appendix, we provide more intuitions about how model capability and sizes affect the complexity of the associations with a set of synthetic experiments over MNIST and multi-layer MLPs.
3.3 Interpreting Example Associations